Circular No. 8206

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2003if IN NGC 1302 Further to IAUC 8201, M. Ganeshalingam and W. Li report the LO-

TOSS/KAIT discovery of a supernova on unfiltered images obtained on Sept. 20.5 (at mag 17.6) and 21.5 UT (at mag 17.7). The new object is located at α = 3^{h}19^{m}52^{s}.61, δ = −26^{o}03′50′′.5 (equinox 2000.0), which is 19′′.3 east and 11′′.9 south of the center of NGC 1302. A KAIT image taken on Jan. 20.2 showed nothing at this position (limiting mag 19.0).

T. Matheson, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2003if, obtained by P. Berlind on Sept. 22.46 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova, ∼ 3 weeks past maximum.

COMET C/2002 T7 (LINEAR) H. Kawakita, Gunma Astrononomical Observatory; J. Watanabe, Na-

tional Astrononomical Observatory of Japan (NAOJ); S. Sasaki and T. Sasaki, University of Tokyo; and N. Takato and T. Fuse, Subaru Telescope, report near-infrared photometry (8′′.4 aperture) of comet C/2002 T7, obtained on Sept. 14.57 UT with the NAOJ 8.2-m Subaru telescope (+ CISCO) at Mauna Kea: J = 12.7, H = 12.3, K = 12.2. The dust coma has a J-band surface brightness of ≈ 20 mag arcsec^{−2} and shows an elongated shape of length 20′′ arcsec in the north-south direction and width 14′′ in the east-west direction. No tail structure was recognized.

V5113 SAGITTARII = NOVA SAGITTARII 2003 No. 2 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

designation V5113 Sgr has been given to this nova (cf. IAUC 8204).

Magnitude estimates for V5113 Sgr forwarded by T. Kato, Kyoto University: Aug. 23.51 UT, [10.6 (T. Hishikura, Yokoyama, Kanagawa, Japan; Tri-X film); Sept. 14.461, [9.9 (Hishikura); 15.443, [10.1 (M. Yamamoto, Aichi, Japan; CCD); 17.424, 10.4 (H. Nishimura, Kakegawa, Shizuoka, Japan; T-Max 400 film); 17.447, 9.7 (K. Haseda, Aichi, Japan; T-Max 400 film; near detection limit); 18.433, 9.1 (Haseda).

Visual magnitude estimates: Sept. 19.934 UT, 9.1 (R. Y. Shida, São Paulo, Brazil); 20.319, 9.2 (J. Bedient, Honolulu, HI); 20.815, 9.2 (P. Schmeer, Bischmisheim, Germany); 21.745, 9.6 (A. Baransky, Pylypovychi, Ukraine); 21.779, 10.0 (K. Hornoch, Lelekovice, Czech Republic).

2003 September 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green